Week of 200127

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Julia (WLCG), Maarten (ALICE + WLCG), Marian (networks), Olga (computing), Petr (ATLAS), Rob (LHCb)
  • remote: Christoph (CMS), Darren (RAL), Dave M (FNAL), David B (IN2P3-CC), Di (TRIUMF), Elena (CNAF), Onno (NLT1), Pepe (PIC), Xin (BNL)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Data reprocessing (carousel) started on Tuesday
    • FTS is not doing well
      • tape systems needs huge number of queued requests for optimal staging
      • FTS doesn't cope well with millions of queued requests
    • CEPH outage - our services heavily relying on CEPH storage
      • our recovery was quite fast but not without manual interventions
      • reliable storage is very important
      • Maarten: fallout from the major network incident (OTG:0054392) in the early hours of Thu

  • CMS reports ( raw view) -
    • No major issues to report
    • CEPH outage from last week was noticed, but did not cause a crisis.

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity:
      • Stripping campaign ongoing
      • Staging will re-start this week based on the stripping activity
    • Issues:
      • no significant issues

Sites / Services round table:

  • ASGC:
  • BNL: NTR
  • CNAF: NTR
  • EGI:
  • FNAL:
    • on Wed there will be a downtime to upgrade the tape dCache to 5.2
      • the disk service is foreseen to be done 2 weeks later
  • IN2P3: NTR
  • JINR: NTR
  • KISTI:
  • KIT: NTR
  • NDGF: All production dCache moved to new HW, bought in the autumn. DBs were moved and upgraded (psql 9.5 -> 12) before Christmas. Old HW empty, will be kept around for a week or two, before repurposing. (Can't participate in phone meeting)
  • NL-T1: The SARA tape system has been down from Saturday 18:00 until this morning.
  • NRC-KI:
  • OSG:
  • PIC: NTR
  • RAL: NTR
  • TRIUMF: For the issues reported in GGUS ticket 144938, the files had been staged in, but due to various reasons, FTS transfers, pin lifetime etc., the files got purged out of tape buffer before transferred to datadisk.

  • CERN computing services: OTG:0054408 - Last week Condor schedds were affected by storage outage.
  • CERN storage services:
  • CERN databases:
  • GGUS:
    • A new release is planned for Wed this week
      • Release notes
      • A downtime has been scheduled for 07:30-10:30 UTC
      • Test alarms will be submitted as usual
  • Monitoring:
    • Two isolated outages during last week:
      • OTG:0054405 - Related with the CEPH outage
      • OTG:0054478 - The physical host holding several InfluxDB databases went into high CPU load and required a manual restart
  • MW Officer:
  • Networks:
    • GGUS:143894 AUSTRALIA-ATLAS - PIC: No recent changes in the network throughput btw. the two sites, performance is degraded, but this is likely due to PIC link capacity
    • GGUS:143770 EELA-UTFSM - UK: Unable to reach UTFSM from UK non-LHCONE sites, reported to JANET/JISC
    • CBPF - CNAF/PIC/IN2P3 - Investigation on-going, CBPF outbound network performance over IPv4 looks good (inbound is problematic), waiting for site to fix perfSONARs to continue
  • Security:

AOB:

Edit | Attach | Watch | Print version | History: r18 < r17 < r16 < r15 < r14 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r18 - 2020-01-27 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback